# Limits of open ASEP stationary measures near a boundary

### v1

@@ Abstract.
Consider the stationary measure of open asymmetric simple exclusion process (ASEP) on the lattice $\{1,\dots,n\}$. Taking $n$ to infinity while fixing the jump rates, this measure converges to a measure on the semi-infinite lattice. In the high and low density phases, we characterize the limiting measure and also show that this convergence occurs in total variation distance on a sublattice of scale $n/\log n$. Our approach involves bounding the total variation distance using generating functions, which are further estimated through a subtle analysis of the atom masses of Askey–Wilson signed measures.
1. Introduction
1.1. Preface
@@
The open asymmetric simple exclusion process (ASEP) serves as a fundamental model for nonequilibrium systems with open boundaries and for Kardar–Parisi–Zhang (KPZ) universality. Over the past five decades, extensive studies have been dedicated to understand the stationary measure of open ASEP, encompassing a wide range of its asymptotic and limiting behaviors. A significant portion of these studies stems from the matrix product ansatz (MPA) method introduced in [1]. This method is notably related to the Askey–Wilson polynomials [2, 3] and processes [4], enabling a rigorous treatment of the phase diagram, large deviations, density profiles and limit fluctuations.
@@
We are interested in a straightforward limit of the open ASEP stationary measures near the boundary. Specifically, we will consider the stationary measure on the lattice $\{1,\dots,n\}$. We fix all the parameters $q,\alpha,\beta,\gamma,\delta$ of the model and take the system size $n$ to infinity. It is known from [5, 6] that such sequences weakly converge, and that the limiting probability measures on $\{0,1\}^{\mathbb{Z}_{+}}$ are (non-unique) stationary measures of certain ASEP systems on the semi-infinite lattice $\mathbb{Z}_{+}$. These types of limits were first studied by Liggett [5] assuming the so-called Liggett's condition, see for example [5, Theorem 1.8 and Theorem 3.10]. Later in Grosskinsky [7, Theorem 3.2] and Sasamoto–Williams [6], a matrix product ansatz was developed to characterize the limiting measures, enabling the studies of their large deviations in Duhart [8] and Duhart–Mörters–Zimmer [9]. The limiting measures were further characterized in Bryc–Wesołowski [4, Theorem 12] in terms of the Askey–Wilson processes, in the `fan region' part of the phase diagram.
@@
In this paper we achieve two main goals. Firstly, using the Askey–Wilson signed measures introduced in a recent work Wang–Wesołowski–Yang [10], we characterize the limiting probability measures on $\{0,1\}^{\mathbb{Z}_{+}}$ in the `shock region' part of the phase diagram. This complements the characterization Bryc–Wesołowski [4, Theorem 12] in the fan region and could serve as a useful tool for further studies of the asymptotics. Secondly, we investigate the following natural question:
@@
In the case $\gamma=\delta=0$ and within the low density phase of the shock region, a recent work by Nestoridi and Schmid [11, Theorem 1.6] provides a growth rate. In this paper, we contribute another (partial) answer to this question: In both the high and low-density phases, the convergence occurs in total variation distance on the leftmost sublattice with a scale of $n/\log n$. We note that in the high density phase, the limiting measures on $\{0,1\}^{\mathbb{Z}_{+}}$ are no longer product Bernoulli measures (as in the low density phase), necessitating a different method.
@@ 1.2. Model and results
Figure 1. Jump rates in the open ASEP.
@@
Figure 2. Phase diagrams for the open ASEP stationary measures. LD, HD, MC respectively stand for the low density, high density and maximal current phases.
@@ 2. $\bullet$
$\mathsf{d}_{TV}(\kappa,\kappa^{\prime}):=\frac{1}{2}\sum_{x\in\{0,1\}^{m}}|\kappa(x)-\kappa^{\prime}(x)|=\max_{A\subseteq\{0,1\}^{m}}\left(\kappa(A)-\kappa^{\prime}(A)\right).$
We equip $\{0,1\}^{\mathbb{Z}_{+}}$ with the infinite product $\sigma$-algebra. The measure $\mu_{n}$ on $\{0,1\}^{n}$ can be regarded as a probability measure on $\{0,1\}^{\mathbb{Z}_{+}}$ by setting $\tau_{i}=0$ for $i\geq n+1$, which we also denote by $\mu_{n}$. As mentioned in the preface, we will characterize the weak limit $\mu_{\infty}$ of measures $\mu_{n}$ as $n\rightarrow\infty$. We will also provide a bound of total variation distance between probability measures $\mu_{n|m}$ and the marginal of $\mu_{\infty}$ on the sublattice $\{0,1\}^{m}$, for $1\leq m\leq n$. This total variation bound will in particular imply that for any sequence $m_{n}$, $n=1,2,\dots$ of scale $n/\log n$, the total variation distance tends to 00.
@@
As a corollary, there exists $s>0$ depending on $A,C$ and $q$ such that for any sequence $\{m_{n}\}_{n=1}^{\infty}$ satisfying $m_{n}\leq s\frac{n}{\log n}$, the total variation distance between $\mu_{n|m_{n}}$ and $\operatorname{Ber}_{m_{n}}\left(\frac{1}{1+C}\right)$ converges to zero as $n\rightarrow\infty$.
The next result is our main theorem in the high density phase. As mentioned in the preface, the limiting measure is no longer product Bernoulli. Instead, the limiting measure is characterized in terms of the Askey–Wilson signed measures introduced in [10].
@@
In the high density phase $A>1$, $A>C$, we assume that $C/A\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $C\geq 1$. Then as $n\rightarrow\infty$, the open ASEP stationary measure $\mu_{n}$ on $\{0,1\}^{n}$ weakly converges to a measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ which we will define as follows:
We define probability measures $\lambda_{m}$ on $\{0,1\}^{m}$ for $m\in\mathbb{Z}_{+}$ by their joint generating functions: For any $t_{1}\leq\dots\leq t_{m}$ within an interval $(1-\varepsilon,1]$ depending on $A,B,C,D$ and $q$,
where $\pi_{t_{1},\dots,t_{m}}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)$ on the RHS above is the multi-dimensional Askey–Wilson signed measure (see Section 2.1 for a brief review). We claim that for $m\in\mathbb{Z}_{+}$, the expression on the RHS above is indeed the generating function of a probability measure $\lambda_{m}$ on $\{0,1\}^{m}$. We also claim that the marginal distribution of $\lambda_{m+1}$ on the leftmost sublattice $\{1,\dots,m\}$ coincides with $\lambda_{m}$. We define the probability measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ by its marginal distribution on the sublattice $\{1,\dots,m\}$ equal to $\lambda_{m}$, for $m=1,2,\dots$.
There exists $H>0$ depending on $A,B,C,D$ and $q$ and $\theta\in(0,1)$ depending on $A,C$ and $q$ such that for any $1\leq m\leq n$ we have:
@@
As a corollary, there exists $s>0$ depending on $A,C$ and $q$ such that for any sequence $\{m_{n}\}_{n=1}^{\infty}$ satisfying $m_{n}\leq s\frac{n}{\log n}$, the total variation distance between $\mu_{n|m_{n}}$ and $\lambda_{m_{n}}$ converges to zero as $n\rightarrow\infty$.
@@
Our total variation distance bounds (1.5) and (1.7) are not expected to be optimal. Hence the growth rate $n/\log n$ of sequences $\{m_{n}\}_{n=1}^{\infty}$ induced by those bounds are also not optimal. It would be an interesting question to ask for the optimal growth rate. Moreover, in the maximal current phase $A<1$, $C<1$, our methods do not provide a tight enough total variation distance bound that induce an interesting growth rate of $\{m_{n}\}_{n=1}^{\infty}$.
1.3. Comparision with a related result
@@
This theorem was proved in [11] using an entirely different method from the present paper. Their method relies on a characterization of the open ASEP stationary measures as convex linear combinations of specific Bernoulli shock measures, under the special condition $ACq^{k}=1$, as shown in Jafarpour–Masharian [14]. When comparing it with Theorem 1.2 in the low density phase, Theorem 1.6 (i.e., [11, Theorem 1.4]) provides a significantly faster growth rate for the sequence $\{m_{n}\}_{n=1}^{\infty}$, albeit within a smaller parameter range, assuming $\gamma=\delta=0$, and within a subregion of the shock region.
1.4. Outline of the paper
Section 2.1 reviews the Askey–Wilson signed measures and open ASEP stationary measures, along with their useful properties. In Section 2.2 we prove the main theorems using technical results that are provided in the three appendices. Appendix A establishes a bound of the total variation distance between two probability measures by their generating functions. Appendix B provides total variation bounds for certain Askey–Wilson signed measures. Appendix C proves a special symmetry of multi-dimensional Askey–Wilson signed measures known as the time reversal.
Acknowledgements
We thank Wlodek Bryc for bringing the paper [11] to our attention, for pointing out an error in the previous version of this manuscript, and for helpful discussions, which, in particular, contributed to Appendix C. We thank Ivan Corwin, Evita Nestoridi, Dominik Schmid, Yizao Wang and Jacek Wesołowski for helpful conversations. The author was partially supported by Ivan Corwin’s NSF grant DMS-1811143 as well as the Fernholz Foundation’s `Summer Minerva Fellows' program.
2. Proofs of the main theorems
We review some backgrounds in Section 2.1 and prove the main theorems in Section 2.2.
2.1. Backgrounds
We first review the definition and some results of Askey–Wilson signed measures following [10]. They were generalized therein from earlier works [15, 4] of Askey–Wilson processes. We then review some useful properties of the open ASEP stationary measures.
@@ 2.1. Backgrounds
Assume $q\in[0,1)$. We denote by $\Omega$ the set of parameters $(a,b,c,d)\in\mathbb{C}^{4}$ satisfying the following three conditions:
@@
    $abcd\notin\{q^{-l}:l\in\mathbb{N}_{0}\}$, where $\mathbb{N}_{0}:=\{0,1,\dots\}$.
@@
$\nu(\mathsf{d}x;a,b,c,d)=f(x;a,b,c,d)\mathds{1}_{|x|<1}+\sum_{x\in F(a,b,c,d)}p(x)\delta_{x},$ (2.1)
@@
$p(y_{k}^{\mathbf{a}})$ $=p_{k}^{\mathbf{a}}(a,b,c,d)=p_{0}^{\mathbf{a}}(a,b,c,d)\frac{q^{k}(1-a^{2}q^{2k})(a^{2},ab,ac,ad)_{k}}{(q)_{k}(1-a^{2})a^{k}\prod_{l=1}^{k}\left((b-q^{l}a)(c-q^{l}a)(d-q^{l}a)\right)},\quad k\geq 1.$ (2.4)
@@
The number $q\in[0,1)$ will be fixed throughout the paper.
@@
For certain parameters $A,B,C,D\in\mathbb{R}$ and on some suitable `time interval' $I\subset\mathbb{R}$ (which will be defined later), we will study the following Askey–Wilson signed measures: For $t\in I$,
@@
Define $U_{t}\subset\mathbb{R}$ to be the (compact) support of $\pi^{(A,B,C,D)}_{t}(\mathsf{d}y)$. We also study Askey–Wilson signed measures: For $s,t\in I$, $s<t$ and $x\in U_{s}$,
@@
For any $t_{1}\leq\dots\leq t_{n}$ in $I$, we define the `multi-dimensional' Askey–Wilson signed measure:
@@
The following result characterizes the open ASEP stationary measures by Askey–Wilson signed measures. This characterization was originally due to [2, 4] and later generalized in [10].
@@
Consider the open ASEP stationary measure $\mu_{n}=\mu_{n}^{(A,B,C,D)}$ on the lattice $\{1,\dots,n\}$. We assume that $ABCD\notin\{q^{-l}:l\in\mathbb{N}_{0}\}$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then there exists $I=[1,1+\varepsilon)$ for some $\varepsilon=\varepsilon(A,B,C,q)>0$ depending on $A,B,C$ and $q$ such that for any $1\leq m\leq n$ and $t_{1}\leq\dots\leq t_{n}$ in $I$, we have:
$\mathbb{E}_{\mu_{n}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{n-m+i}}\right]=\frac{\int_{\mathbb{R}^{m+1}}(2+2x)^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})\pi_{1,t_{1},\dots,t_{m}}\left(\mathsf{d}x,\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)}{\int_{\mathbb{R}}(2+2x)^{n}\pi_{1}(\mathsf{d}x)}.$
@@
This theorem follows from [10, Theorem 1.1] except for the fact that $\varepsilon>0$ can be taken as depending on $A,B,C$ and $q$ but not on $D$. By the arguments in [10], we only need to show that there exists $\varepsilon=\varepsilon(A,B,C,q)>0$ such that for any $s<t$ in $I=[1,1+\varepsilon)$ and $x\in U_{s}$, we have $(a,b,c,d)\in\Omega$, where
@@
Recall that we have assumed $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. One can choose $\varepsilon=\varepsilon(A,B,C,q)>0$ such that $1+\varepsilon<\min\left(1/q,1/B^{2}\right)$ and $1+\varepsilon<1/A^{2}$ if $A<1$, and that the interval $I=[1,1+\varepsilon)$ does not contain elements in $\{Cq^{l}/A:l\in\mathbb{Z}\}$ if $A,C\geq 1$. In the following we will show that for any $s<t$ in $I$ and $x\in U_{s}$, assumption (2) in Definition 2.1 holds for $(a,b,c,d)$ given by (2.8). Note that we always have $b=B\sqrt{t}\in(-1,0]$ for $t\in I$. Also, since $B\sqrt{s},D/\sqrt{s}\in(-1,0]$, any possible atom in $U_{s}$ is generated by either $A\sqrt{s}$ or $C/\sqrt{s}$.
@@
    Let $x\in[-1,1]$ then $c$ and $d$ are complex conjugate pairs with norm $<1$ and (2) holds.

### v2

@@ Abstract
Consider the stationary measure of open asymmetric simple exclusion process (ASEP) on the lattice $\{1,\dots,n\}$. Taking $n$ to infinity while fixing the jump rates, this measure converges to a measure on the semi-infinite lattice. In the high and low density phases, we characterize the limiting measure and provide bounds on the convergence rates in total variation distance. Our approach involves bounding the total variation distance using generating functions, which are further estimated through a subtle analysis of the atom masses of Askey–Wilson signed measures.
82C22,
60K35,
82B23,
Open ASEP,
Askey–Wilson signed measures,
keywords:
[class=MSC]
keywords:
[…]
$0$-0003-2551-7965
1 Introduction
1.1 Preface
@@
The open asymmetric simple exclusion process (ASEP) serves as a fundamental model for nonequilibrium systems with open boundaries and for Kardar–Parisi–Zhang (KPZ) universality. Over the past five decades, extensive studies have been dedicated to understand the stationary measure of open ASEP, encompassing a wide range of its asymptotic and limiting behaviors, including the particle densities [14, 28, 24, 20, 27, 15, 16, 29, 7, 30], limit fluctuations [13, 7, 30], large deviations [15, 16, 9] and open KPZ limits [12, 6, 3, 5]. See survey papers [4, 11] and more references therein. A significant portion of these studies stems from the matrix product ansatz (MPA) method introduced in the seminal work [14]. This method is notably related to the Askey–Wilson polynomials [29, 10] and processes [9].
@@
We are interested in a straightforward limit of the open ASEP stationary measures near the boundary. Specifically, we will consider the stationary measure on the lattice $\{1,\dots,n\}$. We fix all the parameters $q,\alpha,\beta,\gamma,\delta$ of the model and take the system size $n$ to infinity. It is known from [23, 31] that such sequences weakly converge, and that the limiting probability measures on $\{0,1\}^{\mathbb{Z}_{+}}$ are stationary measures of certain ASEP systems on the semi-infinite lattice $\mathbb{Z}_{+}$ with parameters $q,\alpha,\gamma$. We mention that the stationary measures of this semi-infinite ASEP are not unique and are parameterized by the limiting densities at infinity. The aforementioned limits from finite lattices $\{1,\dots,n\}$ to $\mathbb{Z}_{+}$ were first studied by Liggett [23] assuming the so-called Liggett’s condition, see for example [23, Theorem 1.8 and Theorem 3.10]. Later in Grosskinsky [21, Theorem 3.2] and Sasamoto–Williams [31], a matrix product ansatz was developed to characterize the limiting measures, enabling the studies of their large deviations in Duhart [17] and Duhart–Mörters–Zimmer [18]. The limiting measures were further characterized in Bryc–Wesołowski [9, Theorem 12] in terms of the Askey–Wilson processes, in the ‘fan region’ part of the phase diagram.
@@
In this paper we achieve two main goals. Firstly, using the Askey–Wilson signed measures introduced in a recent work Wang–Wesołowski–Yang [30], we characterize the limiting probability measures on $\{0,1\}^{\mathbb{Z}_{+}}$ in the ‘shock region’ part of the phase diagram. This complements the characterization Bryc–Wesołowski [9, Theorem 12] in the fan region and could serve as a useful tool for further studies of the asymptotics. Secondly, we investigate the following natural question:
@@
In the case $\gamma=\delta=0$ and within the low density phase of the shock region, a recent work by Nestoridi and Schmid [25, Theorem 1.4] provides a growth rate. In this paper, we contribute another (partial) answer to this question: In both the high and low-density phases, the convergence occurs in total variation distance on the leftmost sublattice with a scale of $n/\log n$. We note that in the high density phase, the limiting measures on $\{0,1\}^{\mathbb{Z}_{+}}$ are in general no longer product Bernoulli measures (as in the low density phase), necessitating a different method.
@@ 1.2 Model and results
Figure 1: Jump rates in the open ASEP.
@@
Figure 2: Phase diagrams for the open ASEP stationary measures. LD, HD, MC respectively stand for the low density, high density and maximal current phases.
@@
The quantities $\frac{A}{1+A}$ and $\frac{1}{1+C}$ have nice physical interpretations as the ‘effective densities’ near the left and right boundaries of the system, see for example a survey [11, Section 6.2].
@@ 2. $\bullet$
$\mathsf{d}_{TV}(\kappa,\kappa^{\prime}):=\frac{1}{2}\sum_{x\in\{0,1\}^{m}}|\kappa(x)-\kappa^{\prime}(x)|=\max_{A\subseteq\{0,1\}^{m}}|\kappa(A)-\kappa^{\prime}(A)|.$
We equip $\{0,1\}^{\mathbb{Z}_{+}}$ with the infinite product $\sigma$-algebra. The measure $\mu_{n}$ on $\{0,1\}^{n}$ can be regarded as a probability measure on $\{0,1\}^{\mathbb{Z}_{+}}$ by setting $\tau_{i}=0$ for $i\geq n+1$, which we also denote by $\mu_{n}$. As mentioned in the preface, we will characterize the weak limit $\mu_{\infty}$ of measures $\mu_{n}$ as $n\rightarrow\infty$. We will also provide a bound of total variation distance between probability measures $\mu_{n|m}$ and the marginal of $\mu_{\infty}$ on the sublattice $\{0,1\}^{m}$, for $1\leq m\leq n$. This total variation bound will in particular imply that for sequences $m_{n}$, $n=1,2,\dots$ of growth rate $n/\log n$, the total variation distance tends to 00.
@@
As a corollary of the bound (1.5) above, in view of Lemma 2.19, there exists $s>0$ depending on $A,C$ and $q$ such that for any sequence $\{m_{n}\}_{n=1}^{\infty}$ satisfying $m_{n}\leq s\frac{n}{\log n}$, the total variation distance between $\mu_{n|m_{n}}$ and $\operatorname{Ber}_{m_{n}}\left(\frac{1}{1+C}\right)$ converges to zero as $n\rightarrow\infty$.
In the high density phase, as mentioned in the preface, the limiting measure is in general no longer product Bernoulli. Instead, the limiting measure will be characterized in the next definition in terms of the Askey–Wilson signed measures introduced in [30].
Definition 1.3.
In the high density phase $A>1$, $A>C$, we assume that $C/A\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $C\geq 1$. We define probability measures $\lambda_{m}$ on $\{0,1\}^{m}$ for $m\in\mathbb{Z}_{+}$ by their joint generating functions: For some $\varepsilon>0$ and for any $t_{1}\leq\dots\leq t_{m}$ within the interval $(1-\varepsilon,1]$,
where $\pi_{t_{1},\dots,t_{m}}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)$ on the RHS above is the multi-dimensional Askey–Wilson signed measure (see Section 2.1 for a brief review). We will prove in Theorem 1.4 below that there exists $\varepsilon>0$ depending on $A,B,C,D$ and $q$ such that for all $m\in\mathbb{Z}_{+}$, the expression on the RHS above is indeed the generating function of a probability measure $\lambda_{m}$ on $\{0,1\}^{m}$, and that the marginal distribution of $\lambda_{m+1}$ on the leftmost sublattice $\{1,\dots,m\}$ coincides with $\lambda_{m}$. We define the probability measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ by requiring that its marginal distribution on the sublattice $\{1,\dots,m\}$ equals $\lambda_{m}$ for all $m=1,2,\dots$.
The next result is our main theorem in the high density phase.
@@
In the high density phase $A>1$, $A>C$, we assume that $C/A\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $C\geq 1$. Then the probability measures $\lambda_{m}$ on $\{0,1\}^{m}$ for $m\in\mathbb{Z}_{+}$ and the probability measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ in Definition 1.3 are well-defined. Furthermore, as $n\rightarrow\infty$, the open ASEP stationary measure $\mu_{n}$ on $\{0,1\}^{n}$ weakly converges to the measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$.
Moreover, there exists $H>0$ depending on $A,B,C,D$ and $q$ and $\theta\in(0,1)$ depending on $A,C$ and $q$ such that for any $1\leq m\leq n$ we have:
@@
As a corollary of the bound (1.7) above, in view of Lemma 2.19, there exists $s>0$ depending on $A,C$ and $q$ such that for any sequence $\{m_{n}\}_{n=1}^{\infty}$ satisfying $m_{n}\leq s\frac{n}{\log n}$, the total variation distance between $\mu_{n|m_{n}}$ and $\lambda_{m_{n}}$ converges to zero as $n\rightarrow\infty$.
@@
Remark 1.5.
We note that Theorem 1.2 and Theorem 1.4, which respectively concern the low density and high density phases, are not related by the particle-hole duality (Lemma 2.7) of the open ASEP stationary measure. The particle-hole dual of Theorem 1.2 in the low density phase would correspond to a statement in the high density phase, but for the limit as $n\rightarrow\infty$ of a certain transformation of the open ASEP stationary measure—that is, the measure $\widetilde{\mu_{n}}$ on $\{0,1\}^{n}$ defined by $\widetilde{\mu_{n}}(\tau_{1},\dots,\tau_{n})=\mu_{n}(1-\tau_{n},\dots,1-\tau_{1})$ for any $\tau_{1},\dots,\tau_{n}\in\left\{0,1\right\}$, where $\mu_{n}$ is the open ASEP stationary measure. The two limiting measures on $\{0,1\}^{\mathbb{Z}_{+}}$ obtained respectively from $\widetilde{\mu_{n}}$ and $\mu_{n}$ on $\{0,1\}^{n}$ are fundamentally different.
@@
Our total variation distance bounds (1.5) and (1.7) are not expected to be optimal. Hence the growth rate $n/\log n$ of sequences $\{m_{n}\}_{n=1}^{\infty}$ induced by those bounds are also not optimal. It would be an interesting question to ask for the optimal growth rate. Moreover, in the maximal current phase $A<1$, $C<1$, we do not know how to use our methods to obtain a total variation distance bound that is not too loose and that yields a growth rate of $\{m_{n}\}_{n=1}^{\infty}$ that is not too slow. Hence we do not cover this phase in the present paper. It remains an interesting question to find effective total variation distance bounds and the optimal growth rate of $\{m_{n}\}_{n=1}^{\infty}$ in the maximal current phase.
The next result determines exactly when the limiting measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ in the high density phase is a product Bernoulli measure.
Proposition 1.8.
Assume the same conditions as in Definition 1.3, i.e., we are in the high density phase $A>1$, $A>C$, and that $C/A\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $C\geq 1$. Then the measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ introduced in Definition 1.3 is a product Bernoulli measure if and only if $AC=1$, in which case it has density $A/(1+A)$.
Proposition 1.8 will be proved in Section 2.3. We note that when $AC=1$, it is known that the open ASEP stationary measure $\mu_{n}$ on $\{0,1\}^{n}$ is a product Bernoulli measure with density $A/(1+A)$, see for example [19, Appendix A]. Therefore the limiting measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ is also a product Bernoulli measure with the same density.
1.3 Comparision with a related result
@@
This theorem was proved in [25] using an entirely different method from the present paper. Their method relies on a characterization of the open ASEP stationary measures as convex linear combinations of specific Bernoulli shock measures, under the special condition $ACq^{k}=1$, as shown in Jafarpour–Masharian [22]. When comparing it with Theorem 1.2 in the low density phase, Theorem 1.9 (i.e., [25, Theorem 1.4]) provides a significantly faster growth rate for the sequence $\{m_{n}\}_{n=1}^{\infty}$, albeit within a smaller parameter range, assuming $\gamma=\delta=0$, and within a subregion of the shock region.
1.4 Outline of the paper
Section 2.1 reviews the Askey–Wilson signed measures and open ASEP stationary measures, along with their useful properties. In Section 2.2 we prove the main theorems using technical results that are provided in the three appendices. In Section 2.3 we prove Proposition 1.8. Appendix A establishes a bound of the total variation distance between two probability measures by their generating functions. Appendix B provides total variation bounds for certain Askey–Wilson signed measures. Appendix C proves a special symmetry of multi-dimensional Askey–Wilson signed measures known as the time reversal.
{acks}
[Acknowledgments] We thank Wlodek Bryc for bringing the paper [25] to our attention, for pointing out an error in the previous version of this manuscript, and for helpful discussions, which, in particular, contributed to Appendix C. We thank Ivan Corwin, Evita Nestoridi, Dominik Schmid, Yizao Wang and Jacek Wesołowski for helpful conversations. {funding} The author was partially supported by Ivan Corwin’s NSF grants DMS:1811143, DMS:2246576, Simons Foundation Grant 929852, and the Fernholz Foundation’s ‘Summer Minerva Fellows’ program.
2 Proofs of the main theorems
We review some background in Section 2.1 and prove the main theorems in Section 2.2. In Section 2.3 we will prove Proposition 1.8.
2.1 Background
We first review the definition and some results of Askey–Wilson signed measures following [30]. The Askey–Wilson signed measures were introduced in [30] generalizing the Askey–Wilson processes from the earlier works [8, 9]. We later review some useful properties of the open ASEP stationary measures.
@@ 2.1 Background
Assume $q\in[0,1)$. We denote by $\Omega$ the set of parameters $(a,b,c,d)\in\mathbb{C}^{4}$ satisfying the following three assumptions:
@@
    $q^{l}abcd\neq 1$ for all $l\in\mathbb{N}_{0}$, where $\mathbb{N}_{0}:=\{0,1,\dots\}$.
@@
$\nu(\mathsf{d}x;a,b,c,d)=f(x;a,b,c,d)\mathds{1}_{|x|<1}\mathsf{d}x+\sum_{x\in F(a,b,c,d)}p(x)\delta_{x},$ (2.1)
@@
$p(y_{k}^{\mathbf{a}})$ $=p_{k}^{\mathbf{a}}(a,b,c,d)=\frac{p_{0}^{\mathbf{a}}(a,b,c,d)q^{k}(1-a^{2}q^{2k})(a^{2},ab,ac,ad)_{k}}{(q)_{k}(1-a^{2})a^{k}\prod_{l=1}^{k}\left((b-q^{l}a)(c-q^{l}a)(d-q^{l}a)\right)},\quad k\geq 1.$ (2.4)
@@
Parameter $q\in[0,1)$ will be fixed throughout the paper.
@@
For certain parameters $A,B,C,D\in\mathbb{R}$ and on some suitable ‘time interval’ $I\subset\mathbb{R}$ (which will be defined later), we will study the following Askey–Wilson signed measures: For $t\in I$,
@@
Define $U_{t}\subset\mathbb{R}$ to be the (compact) support of $\pi^{(A,B,C,D)}_{t}(\mathsf{d}y)$. We will also study the following Askey–Wilson signed measures: For $s,t\in I$, $s<t$ and $x\in U_{s}$,
@@
For any $t_{1}\leq\dots\leq t_{n}$ in $I$, we define the ‘multi-dimensional’ Askey–Wilson signed measure:
@@
We note that in this paper, the term “Askey–Wilson signed measure” is used to refer to several related objects: the signed measure $\nu(\mathsf{d}x;a,b,c,d)$ defined by (2.1); its parameter specializations $\pi_{t}^{(A,B,C,D)}(\mathsf{d}x)$ and $P_{s,t}^{(A,B)}(x,\mathsf{d}y)$ defined by (2.5) and (2.6) respectively; and the signed measure $\pi^{(A,B,C,D)}_{t_{1},\dots,t_{n}}(\mathsf{d}x_{1},\dots,\mathsf{d}x_{n})$ defined by (2.7), which we sometimes also refer to as the multi-dimensional Askey–Wilson signed measure.
@@
The following result characterizes the open ASEP stationary measures by Askey–Wilson signed measures. This characterization was originally due to [9] in the form of Askey–Wilson processes (see also anterior earlier work [29]) and was later generalized in [30].
@@
Consider the open ASEP stationary measure $\mu_{n}=\mu_{n}^{(A,B,C,D)}$ on the lattice $\{1,\dots,n\}$. We assume that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then there exists $I=[1,1+\varepsilon)$ for some $\varepsilon=\varepsilon(A,B,C,q)>0$ depending on $A,B,C$ and $q$ such that for any $1\leq m<n$ and $t_{1}\leq\dots\leq t_{m}$ in $I$, we have:
$\mathbb{E}_{\mu_{n}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{n-m+i}}\right]\\ =\frac{\int_{\mathbb{R}^{m+1}}(2+2x)^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})\pi_{1,t_{1},\dots,t_{m}}\left(\mathsf{d}x,\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)}{\int_{\mathbb{R}}(2+2x)^{n}\pi_{1}(\mathsf{d}x)}.$
We note that the formula above is the generating function for the open ASEP stationary measure $\mu_{n}$ on $\{0,1\}^{n}$, which involves the multi-dimensional Askey–Wilson signed measure parameterized by $(A,B,C,D)$. In contrast, formula (1.6) in Definition 1.3 gives the generating function for the limiting measure $\lambda_{m}$ on $\{0,1\}^{m}$, which involves the multi-dimensional Askey–Wilson signed measure parameterized by $(A,1/A,C,D)$.
@@
This theorem follows from [30, Theorem 1.1] except for the fact that $\varepsilon>0$ can be taken as depending on $A,B,C$ and $q$ but not on $D$. By the arguments in [30], we only need to show that there exists $\varepsilon=\varepsilon(A,B,C,q)>0$ such that for any $s<t$ in $I=[1,1+\varepsilon)$ and $x\in U_{s}$, we have $(a,b,c,d)\in\Omega$, where
@@
Recall that we have assumed $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. One can choose $\varepsilon=\varepsilon(A,B,C,q)>0$ such that $1+\varepsilon<\min\left(1/q,1/B^{2}\right)$ and $1+\varepsilon<1/A^{2}$ if $A<1$, and that the interval $I=[1,1+\varepsilon)$ does not contain elements in $\{Cq^{l}/A:l\in\mathbb{Z}\}$ if $A,C\geq 1$. In the following we will show that for any $s<t$ in $I$ and $x\in U_{s}$, assumption (2) in Definition 2.1 holds for $(a,b,c,d)$ given by (2.8). Note that we always have $b=B\sqrt{t}\in(-1,0]$ for $t\in I$. Also, since $B\sqrt{s},D/\sqrt{s}\in(-1,0]$, any possible atom in $U_{s}$ is generated by either $A\sqrt{s}$ or $C/\sqrt{s}$. We split the proof into three cases depending on $x\in U_{s}$:
@@
    Let $x\in[-1,1]$ then $c$ and $d$ are complex conjugate pairs with norm $<1$, where the norm of a complex number $z$ refers to its absolute value $|z|$. Hence at most one element in $\{a,b,c,d\}$ has norm $\geq 1$ and assumption (2) vacuously holds.